OpenStudy (anonymous):

Find the real solutions of the equation: 6+√(2x-9)=x Plzz!!! I really need to understand this, THX!!!

OpenStudy (anonymous):

your equation is can be written as: sqrt(2x-9) = x-6 squaring both sides you get: 2x-9 = (x-6)^2 2x-9 = x^2 -12x + 36 FOIL out the right side 0 = x^2 + 14x + 45 set 1 side to zero this is not factorable (easy way) so use the quadratic formula.

OpenStudy (anonymous):

Yup!

OpenStudy (anonymous):

quadratic formula: \[x = \frac {-b \pm \sqrt{b^2-4ac}}{2a}\] where a, b, c are the coefficients of ax^2 + bx + c = 0

OpenStudy (anonymous):

a = 1, b = 14, c = 45. Don't forget to CHECK your answer because you wanted real solutions.

OpenStudy (anonymous):

ok, what will be the solution set, then?? :))

OpenStudy (anonymous):

9 will be the solution??